Consider an economy having following values of Consumption, Investment, Government

 

Spending, and Taxes, 

 

C = 300 + 0.7Yd I = 200 – 200iG = 600T = 100  

 

Derive the aggregate demand equation. Also, calculate the output in an economy


1
Expert's answer
2021-06-03T18:40:22-0400

AD = C + I + G +(X-M)

AD=aggregate demand

C=consumption

I=investment

G=government spending

X=total exports

M=total imports


AD= 300+0.7yd+200-200i+600+100

AD= 0.7yd+200i+1200


EO= Consumption + Investment + Government spending + Taxes

EO= 300+200+600+100= 1200